Joshua is lawyer with a big firm of lawyers.His salary is k78,000 per annum Income tax of k1,050/fortnight is deducted from his pay as well as k100/week for use of a company car.What is his take-home pay each fortnight?
Answer by jorel555(1290)

Annual salary: 78000
78000/26=3000/2 weeks
3000-1050-100=k1850 take home pay every two weeks!!!!!!!!!!!!!!!